OpenStudy (anonymous):

Find the slope of g(θ)=7sinθ-3θ at the point (0,0)

OpenStudy (anonymous):

Hi angelcharm, You have to take the derivative of g with respect to theta. That will give you the gradient (slope) function for any theta. All you need to do then is sub. in your theta value (here, it's 0). So, \[g'(\theta)=7\cos \theta -3 \rightarrow g'(0)=7(0)-3=-3\]

OpenStudy (anonymous):

I'm assuming your point (0,0) is a polar coordinate (i.e. of the form (r, theta)).

OpenStudy (anonymous):

o thanks so much, i got -3 as an answer but i did it a different way but this way is simpler, thanks!

OpenStudy (anonymous):

np become a fan ;)